Using Kirchhoff's rules, find the following. ( 1
Answer #1

I1 = I2 + I3 -----1

I1*R1 + E2 + I2*R2 = E1
I1*2.0k + 61.4 + I2* 3.0k = 70.6 ------2

I1*R1 + E3 + I3*R3 = E1
I1*2.0k + 79.4 + I3*4.0k = 70.6 ------3
Solvinf eq 1,2 & 3

I1 = 0.4 mA
I2 = 2.8 mA
I3 = 2.4 mA


(B)
Potential difference btw c & f,

V = E2 + I2*R2
V = 61.4 + 2.8*3.0
V = 69.8 V
Point c is at higher potential.